The square of an integer is 12 more than the integer. Find integer if 2 or -6 , or -2 or 6 , 3 or -4 or -3 or 4
Answer by unlockmath(1688)

Hello,
Let x be the integer and we now can set up an equation as:
x^2=x+12 Subtract x and 12 from both sides to get:
x^2-x-12=0 Factor this to get:
(x-4)(x+3)=0 Solve for x to result in:
x=4
x=-3
There we go.
Make sense?
RJ
Check out a book at
www.math-unlock.com